# What Do Language Models Hear? Probing for Auditory Representations in Language Models

### v1

@@ Abstract
Jerry Ngo MIT ngop@mit.edu                        Yoon Kim MIT yoonkim@mit.edu
@@ 1 Introduction
Figure 1: (Top) Language (triangle) and sound (circle) representations aligned via Procrustes analysis, visualized via PCA. The language representation is from BERT Devlin et al. (2019) and the audio representation is from PaSST Koutini et al. (2021). The classes are color-coded based on their parent nodes (i.e., human voice, domestic sounds, animal, music) according to the ontology from the FSD50K Fonseca et al. (2021). (Bottom) A zoomed-in portion of the blue region of the top figure, which shows the structural similarities between the language and sound representations for the music category.
Despite being trained only on surface-form strings (i.e., without explicit grounding), language models (LMs) have been shown to learn representations of perceptual concepts that plausibly mirror the grounded, physical representations of those same concepts. Examples of such concepts that have been investigated so far in the literature include color Abdou et al. (2021), direction Patel and Pavlick (2021), size Zhang et al. (2020); Grand et al. (2022), geography Konkol et al. (2017); Liétard et al. (2021); Faisal and Anastasopoulos (2023); Chen et al. (2023), time Gurnee and Tegmark (2023), and even visual representations Ilharco et al. (2021); Merullo et al. (2022); Li et al. (2023). The alignment between an LM’s induced representation of a concept (e.g., the space of word embeddings for colors) and its physical (or human perception-like) representation (e.g., RGB space) has direct implications for how much explicit grounding is necessary for an LM to learn about the “real world” referred to by the textual data on which it was trained. And inasmuch as grounding may be relevant for meaning and understanding, [^1] See Pavlick (2023) and Søgaard (2023) for further discussion on the relationship between grounding and meaning. these findings also have indirect implications for whether LMs can acquire some operationalization of meaning through text-only training Bender and Koller (2020).
@@ 3.1 Models
AudioMAE is a transformer-based model trained as a masked autoencoder on audio spectrograms. AudioMAE is pretrained via self-supervision on the AudioSet dataset Gemmeke et al. (2017), which contains approximately 2 million segments of audio snippets from YouTube along with annotated labels that describe the sound event of the audio snippet (e.g., dog, cat, aircraft, …). [^7] Since the AudioSet dataset contains examples of human speech, AudioMAE’s sound representation is arguably not completely independent of language. The set of possible labels for human speech snippets in the dataset are {male speech, female speech, child speech, conversation, narration, babbling, speech synthesizer}. Self-supervised AudioMAE is trained only on the spectrogram inputs of AudioSet. We also experiment with a supervised, finetuned version of AudioMAE (AudioMAE-FT) that is finetuned as an audio classification model on the same AudioSet dataset.
@@ 4.2 Probing Method
Our primary results make use of a contrastive loss with linear transformations applied to the language/sound representations. We additionally explore two other probes: the Procrustes probe discussed in §2.1 where we learn the matrix $\mathbf{Q}$ only based on the objects in $\mathcal{C}^{\textrm{train}}$; and a non-linear probe where we apply a ReLU non-linearity after projecting when calculating the similarity function. The results are shown Table 2. We generally find that Procrustes probes, which minimize the MSE and additionally constrain the transformations to be orthogonal, underperform the contrastive loss probes. The linear probe outperforms the non-linear probe.
@@ 5 Discussion and Limitations
This work only explored whether sound representations that were learned by an auxiliary audio processing model were encoded through text. Here we found that sound representations that are more likely to encode auditory (i.e., human perception-like) information were more aligned to the text representations than sound representations from purely self-supervised models which were just trained on spectrograms, and thus more likely to encode acoustic information. However, even the self-supervised audio models implicitly encode human perception-aligned priors given that the input data consisted of snippets of audio that corresponded to different sound events, which itself is derived from humans. It would be interesting to see whether it is possible to probe out even more low-level representations of objects (e.g., raw spectrograms, pixels) from language models.
@@ 6 Related Work
Language models have been shown to encode much linguistic information in their contextualized representations (Tenney et al., 2019; Liu et al., 2019; Jawahar et al., 2019) and attention distributions (Clark et al., 2019; Vig and Belinkov, 2019). Building on top of these more linguistically-oriented probes, there has been mounting recent evidence that language models trained on just text are able to meaningfully encode a surprising amount of grounded or extralinguistic information, such as color Abdou et al. (2021), direction Patel and Pavlick (2021), size Zhang et al. (2020); Grand et al. (2022), geography Konkol et al. (2017); Liétard et al. (2021); Faisal and Anastasopoulos (2023); Chen et al. (2023), time Gurnee and Tegmark (2023), visual representations (Ilharco et al., 2021; Merullo et al., 2022; Li et al., 2023), character-level information of word-pieces (Kaushal and Mahowald, 2022), and representations of meaning (Li et al., 2021). Our work extends this line of to sounds and investigates the extent to which language models trained on text-only can encode auditory representations.
@@ Appendix A Appendix
Table 3 shows the full numeric results for our main probing experiments.

### v2

@@ Abstract
Jerry Ngo               Yoon Kim Massachusetts Institute of Technology ngop@mit.edu    yoonkim@mit.edu
@@ 1 Introduction
Figure 1: (Top) Language (triangle) and sound (circle) representations aligned via Procrustes analysis Schönemann (1966), visualized via PCA. The language representation is from BERT Devlin et al. (2019) and the audio representation is from PaSST Koutini et al. (2021). The classes are color-coded based on their parent nodes (i.e., human voice, domestic sounds, animal, music) according to the ontology from the FSD50K Fonseca et al. (2021). (Bottom) A zoomed-in portion of the blue region of the top figure, which shows the structural similarities between the language and sound representations for the music category.
Despite being trained only on surface-form strings (i.e., without explicit grounding), language models (LMs) have been shown to learn representations of perceptual concepts that plausibly mirror the grounded, physical representations of those same concepts. Examples of such concepts that have been investigated so far in the literature include color Abdou et al. (2021), direction Patel and Pavlick (2021), size Zhang et al. (2020); Grand et al. (2022), geography Konkol et al. (2017); Liétard et al. (2021); Faisal and Anastasopoulos (2023); Chen et al. (2023), time Gurnee and Tegmark (2023), and even visual representations Ilharco et al. (2021); Merullo et al. (2022); Li et al. (2023). The alignment between an LM’s induced representation of a concept (e.g., the space of word embeddings for colors) and its physical (or human perception-like) representation (e.g., RGB space) has direct implications for how much explicit grounding is necessary for an LM to learn about the “real world” referred to by the textual data on which it was trained. And inasmuch as grounding may be relevant for meaning and understanding, [^1] See Pavlick (2023) and Søgaard (2023) for further discussion on the relationship between grounding and meaning. these findings also have indirect implications for whether LMs can acquire (some operationalization of) meaning through text-only training Bender and Koller (2020).
@@ 3.1 Models
AudioMAE is a transformer-based model trained as a masked autoencoder on audio spectrograms. AudioMAE is pretrained via self-supervision on the AudioSet dataset Gemmeke et al. (2017), which contains approximately 2 million segments of audio snippets from YouTube along with annotated labels that describe the sound event of the audio snippet (e.g., dog, cat, aircraft, …).777 Since the AudioSet dataset contains examples of human speech, AudioMAE’s sound representation is arguably not completely independent of language. The set of possible labels for human speech snippets in the dataset are {male speech, female speech, child speech, conversation, narration, babbling, speech synthesizer}. Self-supervised AudioMAE is trained only on the spectrogram inputs of AudioSet. We also experiment with a supervised, finetuned version of AudioMAE (AudioMAE-FT) that is finetuned as an audio classification model on the same AudioSet dataset.
@@ 4.2 Probing Method
Our primary results make use of a contrastive loss with linear transformations applied to the language/sound representations. We additionally explore two other probes: the Procrustes probe discussed in §2.1 where we learn the matrix $\mathbf{Q}$ only based on the objects in $\mathcal{C}^{\textrm{train}}$; and a non-linear probe where we apply a ReLU non-linearity after projecting when calculating the similarity function.
The results are shown Table 2. We generally find that Procrustes probes, which minimize the MSE and additionally constrain the transformations to be orthogonal, underperform the contrastive loss probes. The linear probe outperforms the non-linear probe.
@@ 5 Discussion and Limitations
This work only explored whether sound representations that were learned by an auxiliary audio processing model were encoded through text. Here we found that sound representations that are more likely to encode auditory (i.e., human perception-like) information were more aligned to the text representations than sound representations from purely self-supervised models which were just trained on spectrograms, and thus more likely to encode acoustic information. However, even the self-supervised audio models implicitly encode human perception-aligned priors given that the input data consisted of snippets of audio that corresponded to different sound events, which itself is derived from humans. It would be interesting to see whether it is possible to probe out even more low-level representations of objects (e.g., raw spectrograms, pixels) from language models. Similarly, as discussed in footnote 7 our audio representations are not completely independent of language as their training sets included a significant amount of human speech. It would therefore be interesting to see if audio models trained without any human speech learn representatons that can be aligned to language models.
@@ 6 Related Work
Language models have been shown to encode much linguistic information in their contextualized representations (Tenney et al., 2019; Liu et al., 2019; Jawahar et al., 2019) and attention distributions (Clark et al., 2019; Vig and Belinkov, 2019). Building on top of these more linguistically-oriented probes, there has been mounting recent evidence that language models trained on just text are able to meaningfully encode a surprising amount of grounded or extralinguistic information, such as color Abdou et al. (2021), direction Patel and Pavlick (2021), size Zhang et al. (2020); Grand et al. (2022), geography Konkol et al. (2017); Liétard et al. (2021); Faisal and Anastasopoulos (2023); Chen et al. (2023), time Gurnee and Tegmark (2023), visual representations (Ilharco et al., 2021; Merullo et al., 2022; Li et al., 2023), character-level information of word-pieces (Kaushal and Mahowald, 2022), and representations of meaning (Li et al., 2021). LM-derived similiarity measures have also been shown to correlate with human-derived similarity measures across a number of modalities Marjieh et al. (2023a, b). The present work extends the probing-based line of work to sounds, and investigates the extent to which language models trained on text-only can encode auditory representations.
@@ Appendix A Appendix
We show the full numeric results for our main probing experiments in table 3 and for classes accuracy in figure 7.
@@
(a) AudioMAE
(b) AudioMAE-FT
(a) PaSST
(b) PANN
Figure 7: All class accuracies for different audio models. Each class’s accuracy is averaged across 5 train/test sets and different language models.